OpenStudy (anonymous):

Express as a single logarithm. log6496 – log62 I know logs pretty well but I don't know how to express 2 logs as a single log, can someone please help??

OpenStudy (anonymous):

I meant log base 6 of 496 - log base 6 of 2

OpenStudy (anonymous):

\[\log_{6} 496-\log_{6} 2=\log_{6} \frac{ 496 }{ 2 }=\log_{6} 248\]

OpenStudy (anonymous):

That's what I thought thank you! What about, log base 5 of 50 + log base 5 of 2.5?

OpenStudy (anonymous):

I know it has to equal something in the final answer and that confuses me

Nnesha (nnesha):

quotient rule\[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\] you should familiar with these

Nnesha (nnesha):

there is a plus sign so which property you should apply ?

OpenStudy (anonymous):

the product property

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

I am doing a test review and all of the multiple choice answers for the question equal something, like one answer is, log base 5 of 125=3

OpenStudy (anonymous):

How do you find what it equals?

Nnesha (nnesha):

log base 5 of 50 + log base 5 of 2.5 for this question ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

\[\huge\rm log_5 50+ \log_5 2.5\] is this ur question ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

there isn't an equal sign..

OpenStudy (anonymous):

These are the possible answers: \[\log_{5}125=3 \] \[\log_{5}112.5=2.93 \] \[\log_{5}112.5=22.5 \] \[\log_{5}125=625 \]

Nnesha (nnesha):

alright then you have to apply the `change of base` formula \[\huge\rm log_b a= \frac{ \log a }{ \log b }\]

Nnesha (nnesha):

\[\huge\rm log_5 50+ \log_5 2.5\] first) how would you write this in single log form b applying product property ?

OpenStudy (anonymous):

\[\log_{5}125 \] ?

Nnesha (nnesha):

yes right now use change of base formula

OpenStudy (anonymous):

How do I know what to change the base to?

Nnesha (nnesha):

base would stay the same hmm what do you mean ? o.O here is an example \[\log_2 3 =\frac{ \log(3) }{ \log(2) }\]

OpenStudy (anonymous):

oh okay! I was looking at my notes and I read the equation wrong!

Nnesha (nnesha):

ohh

OpenStudy (anonymous):

soooo is the answer \[\log_{5}125=3 \]

Nnesha (nnesha):

if log_5 (125) is equal to 3 then yes did you use the formula ?

Nnesha (nnesha):

how would you use `change of base formula` for \[\large\rm \log_5 125\]

OpenStudy (anonymous):

\[\frac{ \log(125) }{ \log(5) }\]

Nnesha (nnesha):

perfect!

OpenStudy (anonymous):

Awesome! This makes sense now! Thank you so much! You are very smart

Nnesha (nnesha):

glad to hear that & thanks & good job! good luck !

OpenStudy (anonymous):

Thanks! :)